hope someone could help me with this
Suppose a farmer has 160 feet of fencing to make a rectangular barnyard enclosure. If the enclosure is x feet long, express the area A of the enclosure as a function of the length. Give an answer by describing your solution.
The solution is one of the following:

1. A(x) = x^2 – 40x
2. B. A(x)= square root 25,600-x^2
3. A(x) = 160x – x^2
4. A(x) = 80x – x^2

Which one do you think is correct? show your work

Answer by nerdybill(7384)

Since, perimeter is:
L+L+W+W
or
2(L+W)
The problem GAVE us the length as x.
2(x+W)
.
Since the problem told us that the total length of fencing is 160 feet we have:
160 = 2(x+W)
Solving for W:
80 = x+W
80 - x = W
.
Well, we all know that the area of a rectangle is length*width we get:
x(80 - x)
= 80x - x^2